OpenStudy (anonymous):

simplify each exponential expression x^2 * x^5

OpenStudy (thewayiam983):

x^2*x^5=x^(2+5)=x^7 or (x*x)*(x*x*x*x*x)=(x*x*x*x*x*x*x) which is x times itself seven times